A mass of 0.4kg is rotated by a string at constant speed v in a vertical circle. R=1m. If the minimum tension is 3N. Calculate

a) v

b) maximum tension when string is horizontal


1
Expert's answer
2021-08-03T11:41:14-0400

"m =0.4kg;R = 1m;T_{min}= 3N"

"\vec g = 9.8\frac{m}{S^2}"

"\vec F = m\vec a"

"\vec a = \frac{v^2}{R}= v^2\text{ etc. }R=1"

"\vec F=\vec F_g+\vec{T}"

"\vec F_g = m\vec g"

"\vec T= \vec F- \vec F_g= m(\vec a- \vec g)"

"T_{min} = m (a-g)"

"a = \frac{T_{min}}{m}+g= \frac{3}{0.4}+9.8=17.3\frac{m}{s^2}"


"a) a = \frac{v^2}{R}"

"v= \sqrt{aR}= \sqrt{17.3*1}=4.16\frac{m}{s}"


"b)\text{ maximum tension when string is horizontal}"

"\vec F=\vec F_g+\vec{T}"

"\text{Projection of forces on the horizontal axis:}"

"F_x=T"

"T = |\vec F|= ma= 0.4*17.3= 6.92N"


"\text{Answer: }a)v=4.16\frac{m}{s};b)T = 6.92 N"